OpenStudy (anonymous):

An atom has an atomic number of 19 its atomic mass is 39 how many neutrons does it have?

OpenStudy (unklerhaukus):

the atomic number is the number of protons the mass number is is the sum of the protons and neutrons

OpenStudy (anonymous):

20. 39-19=20

OpenStudy (anonymous):

Mass = Protons + Neutrons Neutorns = Mass - Protons 39 - 19 = 20

OpenStudy (anonymous):

Atomic Mass = protons + neutrons 39 = 19 + neutrons neutrons = 39 - 19 = 20